Q.3 Answer all parts Total 20 marks] The Engineering Manager in A Company Ltd. is responsible for maximising the profit from four product lines A, B, C and D. They have been provided with the following information; Direct materials (kgs) 80 95 70 Direct labour (hrs) Welding Assembly Finishing 4.5 9 4 4.5 9.5 4 3.5 4 8 4 Selling price per unit (E) 1500 2500 3000 1000 280 Max.sales for the period in units) 400 300 450 250 250 Budgeted production (in units) 250 250 Additional information is also provided i. There are no closing stocks ii. Material costs are estimated at £5.00/kg ii. The labour rate for the assembly and finishing operations is £15.00/hr and £20/hr for the welding operation iv. The skilled labour required for the welding process is in short supply although there is just enough to meet the current budgeted production (a) Prepare a statement showing the contribution from each product based on the budgeted production (b) Calculate the contribution per unit of limiting factor (c) Prepare a statement showing the maximum total contribution that can be (8 marks) (2 marks) obtained. Explain your rationale (10 marks)
Answer #1

Answer (a):

A Company Limited Contribution from each product based budgeted production Total Per Unit Total (250 units) Per Unit Total (2

Direct material cost per unit = Direct material per unit in kgs * £5 per kg

Direct labor cost per unit (for welding) = Direct labor required for welding per unit * £20

Direct labor cost per unit (for assembly and finishing) = Direct labor hours required for assembly or finishing per unit * £15

Answer (b):

Limiting factor is Skilled labor required for welding process.

Contribution per unit of limiting factor is calculated below:

A Company Limited Contribution from each product based budgeted production (in £) Per Unit lPer Unitl Per Unitl Per Unit 1,500| 2,500| 3,000| 1,000 Sale value Less, Variable cost: Direct Material 400.00 450.00 475.00 350.00 Direct labor Weldin Assembl Finishin Total Variable cost 80.00 90.00 90.00 70.00 120.00 135.00 142.50 90.00 60.00 60.00 60.00 45.00 660.00 735.00 767.50 555.00 840.00 1,765.00| 2,232.50445.00 Contribution Direct labor(Hours) for Welding per unit Contribution per unit of limiting factor 4 4.5 210.00 392.22 496.11 127.14 Preference 2 4

Answer (c):

Skilled labor is in short supply but they are just enough to meet budgeted production.

Budgeted production is 250 units of each product

Available skilled labor = 250 * (4 + 4.5 + 4.5 + 3.5) = 4125 hours

As calculated in answer b above:

The highest contribution per limiting factor is from product C:

Hence to maximize contribution:

1. Produce Product C up to units as Max Sales

Skilled labor required = 280 * 4.5 = 1260 hours

Balance of skilled labor = 4125 - 1260 = 2865 hours

2. Next best contribution per unit of limiting factor is from Product B

Produce Product B up to units as Max Sales

Skilled labor required = 300 * 4.5 = 1350 hours

Balance of skilled labor = 2865 - 1350 =1515 hours

3. Next best contribution per unit of limiting factor is from Product A

Maximum sales units of product A = 400 units

Available skilled labor = 1515 hours

Produce product A = 1515/ 4 = 378.75 units or 379 units

As such optimized production plan will be:

Product A = 378.75 units (for calculations we are taking units in fraction)

Product B = 300 units

Product C = 280 units

Product D = 0 unit

Maximum total contribution that can be obtained = £1,472,750

The statement is as follows:

A Company Limited Contribution from each product based budgeted production Total Per UnitTotal (378.75 units) Per Unit Total (300 units)Per UnitTotal (280 units) Sale value Less, Variable cost: 1,500.00 568,125.002,500 750,000.00 3,000 840,000.00 2,158,125.00 Direct Material 400.00 151,500.00 450.00 135,000.00 475.00 133,000.00 419,500.00 Direct labor Weldin Assembl Finishin Total Variable cost 80.00 120.00 60.00 660.00 840.00 30,300.00 90.00 45,450.00 135.00 22,725.00 60.00 249,975.00 735.00 318,150.001,765.00 27,000.00 90.00 40,500.00 142.50 18,000.00 60.00 220,500.00 767.50 529,500.00| 2,232.50 82,500.00 39,900.00 125,850.00 57,525.00 214,900.00 685,375.00 625,100.00 1,472,750.00 25,200.00 16,800.00 Contribution